Simplify the following expression

Cos ( x+ Pi over 3) - sin ( x + Pi over 6 )
1
Expert's answer
2020-10-26T12:57:16-0400

"cos(x+\frac{\pi}{3}) - sin(x+\frac{\pi}{6})"

"cos(x+\frac{\pi}{3})= cosxcos\frac{\pi}{3} - sinxsin\frac{\pi}{3}"

"sin(x+\frac{\pi}{6})= sinxcos\frac{\pi}{6}+cosxsin\frac{\pi}{6}"

"cos(x+\frac{\pi}{3}) - sin(x+\frac{\pi}{6})=" "cosxcos\frac{\pi}{3} - sinxsin\frac{\pi}{3}" "-sinxcos\frac{\pi}{6}-cosxsin\frac{\pi}{6}"

therefore

"cosxcos\frac{\pi}{3}-sinxsin\frac{\pi}{3}-sinxcos\frac{\pi}{6}-cosxsin\frac{\pi}{6}"

this becomes

"\frac{1}{2}cosx-\frac{1}{2}cosx-\frac{\sqrt3}{2}sinx-\frac{\sqrt3}{2}sinx"

"0-\sqrt3sinx = -\sqrt3 sinx"